OpenStudy (ars22):

how do you solve |2k-3|=3

OpenStudy (anonymous):

|2k-3|=3 which mean (2k-3 ) could be either 3 or -3

OpenStudy (anonymous):

(2k-3)=-3 2k=0 k=0

OpenStudy (anonymous):

2k-3=3 2k=6 k=3

OpenStudy (lilg132):

2k - 3 = 3 or 2k - 3 = -3 2k = 3 + 3 2k = 6 k = 3 or 2k - 3 = -3 2k = 0 k = 0